OpenStudy (anonymous):

integration of sin^6x+sin^7x

OpenStudy (anonymous):

quick question

OpenStudy (anonymous):

what does the symbol next to the sin mean?

OpenStudy (anonymous):

multiplication?

OpenStudy (anonymous):

any1?

OpenStudy (owlfred):

\[\int (\sin^6 x + \sin ^7 x) dx\]

OpenStudy (anonymous):

i see now thnx!

OpenStudy (anonymous):

okay i have no clue what sin means so i will just help another question.. I haven't gotten there yet

OpenStudy (owlfred):

@ganeshie8 will certainly help you out on this one.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aum):

https://www.youtube.com/watch?v=RjvYKVmV2EQ